# Supersonic shear waves in dipolar supersolids

### v2

@@ Abstract
Dipolar supersolids—quantum states which are simultaneously superfluid and solid—have had their superfluid nature rigorously tested, while its solid nature remains uncharted. Arguably, the defining characteristic of a solid is the existence of elastic shear waves. In this work, we investigate transverse wave packet propagation in dipolar supersolids with triangular and honeycomb structure. Remarkably, the honeycomb supersolid displays anomalous dispersion, supporting waves traveling faster than the transverse speed of sound: a supersonic shear wave. For both supersolid phases, we calculate the shear modulus, a key parameter that quantifies the material’s rigidity. Our findings are pertinent to current experimental efforts scrutinizing the fundamental properties of supersolids.
@@
One other such quantum phase that should support shear waves is the supersolid. A supersolid is a fascinating and counterintuitive state of matter that combines the properties of superfluidity with crystalline order [7, 8, 9, 10, 11, 12]. In this exotic phase, a material exhibits both the rigidity and periodic structure characteristic of an ordered solid, and the inviscid flow typically associated with superfluids. This duality emerges from multiple spontaneously broken symmetries: a global U(1) phase symmetry, and both translation and rotation symmetries [13]. Ultracold atoms have now established themselves as the leading platform to study supersolidity, with varied approaches including cavity settings [14, 15, 16], atoms with spin-orbit coupling [17], and dipolar atoms exhibiting intrinsic long-range interactions [18, 19, 20, 21]. Understanding this phase in the context of ultracold matter may unveil new behavior in solid state materials [22], helium systems [23, 24, 25, 26], and neutron stars [27, 28, 29].
@@
Here, we propose a protocol to excite shear waves in two-dimensional dipolar supersolids, and measure their resulting propagation velocities. We investigate this protocol for triangular and honeycomb lattice structures, which are ground state solutions for a system of ultracold dipolar atoms under planar confinement [40, 41, 42]. Our results reveal distinct dispersive phenomena between the triangular and honeycomb phases, which we probe through the current density dynamic structure factor (DSF). These investigations show that the honeycomb structure can support a refractive index smaller than unity, thus supporting shear waves traveling faster than the transverse speed of sound. This phenomenon is analogous to the anomalous phonon dispersion relation observed in superfluid helium-4 [43, 44, 45, 23]. Finally, we find the relationship between the wave packet propagation velocities and the transverse speed of sound, and estimate the shear modulus of dipolar supersolids.
@@
This full system is well described by a beyond mean-field theory, known as the extended Gross-Pitaevskii equation (eGPE) [47, 48, 49, 50]. This governs the three-dimensional macroscopic wavefunction $\Psi(\textbf{r},t)$, and takes into account single-particle dynamics, two-body interactions, and quantum fluctuations. Due to the azimuthal symmetry of the interaction, dipolar supersolids break at most two translational symmetries, giving rise to triangular, stripe, or honeycomb two-dimensional (2D) lattice structures [40, 41, 42]. Thus, following Ref. [41], we are able to reduce the complexity of this model to an infinite planar geometry, solving for the dimensionally reduced wave function $\psi(\bm{\rho},t)$ with variational width $\sigma$. We fix the areal number density $n_{\text{2D}}$, and perpendicular harmonic trap (in $z$) with frequency $\omega_{z}=2\pi\times 72.4\mathrm{\,Hz}$. This model and our numerical methods are detailed in the Appendix.
@@
Shear waves in triangular supersolids—In Fig. 1(a), we show the initial perturbation of a single row of the crystal away from its equilibrium position, which adds elastic tension to the lattice. We shift the central row by $\Delta L=0.125L$, where $L$ is the lattice spacing  [^2] We have ensured that the observed shear wave propagation does not vary with the magnitude of this shift for smaller perturbations.. Evolving the system in real time, we observe a longitudinally propagating wave packet (along $\hat{y}$), leaving behind transverse crystal oscillations (along $\hat{x}$). We then track the deviation of the center of mass (COM) of each row from its equilibrium position, as shown in Fig. 1(b). We track both the peak and edge of this wave packet, defining the latter as a COM deviation larger than 5% of the peak amplitude. The peak velocity is non-linear, starting at $\approx 0.22\,$mm/s and increases with the propagation time to $\approx 0.24\,$mm/s after 750 ms, and the edge velocity always exceeds the peak velocity at $\approx 0.26\,$mm/s, due to the dispersive nature of the wave packet. The existence of shear waves implies a non-zero shear modulus of dipolar supersolids, and proves the solid-like nature of the state, a key result of this work.
@@
Figure 1(c) shows half of the first Brillouin Zone (BZ), $0<k_{y}<k_{\text{BZ}}=\pi/L$, of the current density DSF, revealing the three gapless (Goldstone) modes inherent to a 2D supersolid [59]. As apparent from the large amplitude positive $\mathcal{S}_{j}$, the response of the system is dominated by $j_{x}$ fluctuations, revealing the transverse branch $\omega_{t}(k_{y})$. Only weak longitudinal crystal and superfluid phonons are induced, as expected from the dislocation in the density at the boundary of the initial perturbation. On the right-hand side of (c), we also highlight the dominant current density fluctuation by taking $\mathcal{F}_{t}[j_{x}(\bm{\rho},t)]=j_{x}(\bm{\rho},\omega)$, and isolating the mode with frequency $\omega_{t}(k_{\text{BZ}})=\omega_{\text{BZ}}$ [60]. We note that $\omega_{\text{BZ}}$ is equal to the oscillation frequency of the central row COM ($y=0$), shown in (b). This mode corresponds to a transverse phonon with wave number $k_{\text{BZ}}$, i.e. a fluctuation where the current alternates between each row.
We can extract three characteristic velocities from the transverse branch: 1) the transverse speed of sound $c_{t}=\lim_{k_{y}\rightarrow 0}\omega_{t}/k_{y}$, 2) the phase velocity $v_{p}(k_{y})=\omega_{t}/k_{y}$, which defines the velocity of a wave with given $k_{y}$, and 3) the group velocity, $v_{g}(k_{y})=\partial\omega_{t}/\partial k_{y}$, which defines the velocity of the wave packet envelope and thus the velocity of energy propagation in the medium  [^3] We extract these quantities from the current density DSF by fitting the discrete transverse branch with a fourth-order polynomial.. For the triangular supersolid, both $v_{g}$ and $v_{p}$ monotonically decrease from $c_{t}$ with increasing wave number, and subsequently the refractive index, $n=c_{t}/v_{p}$, increases from unity with $k_{y}$. This behavior is referred to as normal dispersion and is a hallmark of conventional crystalline solids [62].
@@
We can pinpoint the origin of this unexpected finding from the current density DSF and the corresponding characteristic velocities, as shown in Figs. 2(c)-(d)  [^4] Note that as the perturbation dislocates the high density region, we excite many branches, but still dominantly transverse crystal modes.. The transverse branch displays a subtle upward concavity for intermediate wave vector ($0.2\lesssim k_{y}L/\pi\lesssim 0.4$), supporting supersonic $v_{g}$ and $v_{p}$ ($>c_{t}$), and a sub-unity refractive index. This is known as anomalous dispersion and has significant implications for transverse phonon decay. The superlinear dispersion allows for Beliaev damping of transverse phonons, where one phonon spontaneously decays into two [64, 65]. Though rare, similar phenomena has been observed in other systems such as crystalline materials [66, 67], and superfluid helium-4 [44], the latter being the subject of thirty years of intense experimental and theoretical debate [23]. To our knowledge, this constitutes the first quantum solid to display anomalous dispersion, the main result of this Letter.
@@
Figure 3: Shear wave velocities and the shear modulus for triangular $a_{s}/a_{\text{dd}}=0.75$ and honeycomb $a_{s}/a_{\text{dd}}=0.775$ supersolids. (a) Points indicate the estimated transverse speed of sound, $c_{t}$, with 95% confidence error bar from a linear fit of the transverse branch at low $k$. Shaded region is bounded by the slowest peak COM velocity (dashed curve) and edge COM velocity (dotted curve). (b) Superfluid fraction. (c) Shear modulus, $\mu$, calculated with the corresponding velocities in (a).
Figure 3 shows the required ingredients in order to estimate the shear modulus from our measured time-dependent velocities $v_{s}$. We generalize our results to a wide-range of 2D densities, whilst maintaining the ground state crystal structure. The shaded region in (a) demarcates the range of velocities that could be observed in an experiment. The lower bound is given by the initial peak velocity (which increases over time), and the upper bound is the edge velocity. Combined with the superfluid fraction calculation of (b), we show that the shear moduli of both states are on the order of fN/mm  [^6] As the crystal is 2D, the units are in Newtons per length.. Thus, the small–but non-zero–values of the shear moduli definitively show that the dipolar supersolid phases are solid  [^7] We have verified for both states that these observations are only weakly dependent on the scattering length..
Consider the experimental implementation of this protocol for the triangular supersolid. The edge velocity matches closely with our extracted $c_{t}$ [Fig. 3(a1)], however due to fluctuations such as temperature effects, tracking the edge of the shear wave may be below the fidelity of experiments. Furthermore, in finite size systems the propagation velocity can be greatly reduced at the boundary of the system, as we show in the Appendix. Thus, experimental estimates may drastically underestimate the shear modulus. At much larger densities than shown here, the transverse branch becomes soft, and the shear wave velocity tends to zero. However, this is in the parameter regime where the stripe phase is the ground state solution [41].
@@
Acknowledgements—We acknowledge invaluable conversations with F. Ferlaino, R. Bisset, E. Poli, P. B. Blakie, N. Liebster, M. Olimpo, and the Innsbruck dipolar team. We are particularly grateful to P. B. Blakie for sending us data from Refs. [41] and [71] to benchmark our results. We acknowledge financial support by the ESQ Discovery programme (Erwin Schrödinger Center for Quantum Science & Technology), hosted by the Austrian Academy of Sciences (ÖAW), and the European Research Council through the Advanced Grant DyMETEr (10.3030/101054500), a NextGeneration EU Grant AQuSIM through the Austrian Research Promotion Agency (FFG) (No. FO999896041). T.B. acknowledges funding from FWF Grant No. I4426.
Appendix
Appendix A—To model our infinite planar system, the 3D wavefunction can be separated into $\Psi(\textbf{r},t)\equiv\Psi(\bm{\rho},z,t)=\psi(\bm{\rho},t)\chi_{\sigma}(z)$, where $\bm{\rho}=(x,y)$ and $\chi_{\sigma}(z)=\pi^{-1/4}\sigma^{-1/2}e^{-z^{2}/2\sigma^{2}}$ with variational width $\sigma$. Following Ref. [41], we substitute this ansatz into the 3D eGPE energy functional and perform the integration over $z$ to find the new total energy functional $E=E_{\perp}+E_{z}$, where
$E_{\perp}=\int\text{d}^{2}\bm{\rho}\,\psi^{*}\bigg{[}-\frac{\hbar^{2}\nabla^{2}}{2m}+\frac{1}{2}\Phi+\frac{2}{5}g_{\text{QF}}|\psi|^{3}\bigg{]}\psi\,,$ (5)
and $E_{z}=\hbar\omega_{z}(l_{z}^{2}/4\sigma^{2}+\sigma^{2}/4l_{z}^{2})$, where $l_{z}=\sqrt{\hbar/m\omega_{z}}$ is the harmonic oscillator length. The two-dimensional interaction potential can be evaluated in Fourier space as
$\Phi(\bm{\rho},t)=\mathcal{F}^{-1}\left[\tilde{U}(\bm{k})\mathcal{F}\left[|\psi(\bm{\rho},t)|^{2}\right]\right]\,,$ (6)
with $\bm{k}=(k_{x},k_{y})$, (inverse) Fourier transform $\mathcal{F}$ ($\mathcal{F}^{-1}$), and
$\tilde{U}(\bm{k})=\frac{\sqrt{2\pi}\hbar^{2}}{\sigma m}\left(a_{s}+a_{\text{dd}}\left[2-3G_{0}\left(\sigma\bm{k}/\sqrt{2}\right)\right]\right)\,,$ (7)
and $G_{0}(q)=\sqrt{\pi}qe^{q^{2}}\text{erfc}(q)$, is the dimensionally reduced Eq. (1). Finally, we have also introduced the following scaled quantum fluctuation term [87]
$g_{\text{QF}}=\frac{128\hbar^{2}}{3m}\sqrt{\frac{2a_{s}^{5}}{5\pi^{1/2}\sigma^{3}}}\,\,\left[1+\frac{3}{2}\left(\frac{a_{\mathrm{dd}}}{a_{s}}\right)^{2}\right]\,.$ (8)
One can calculate $\delta E_{\perp}/\delta\psi^{*}$ to obtain the quasi-2D variational eGPE
$i\hbar\frac{\partial\psi}{\partial t}=\bigg{[}-\frac{\hbar^{2}\nabla^{2}}{2m}+\Phi_{\sigma}+g_{\text{QF}}|\psi|^{3}\bigg{]}\psi\,.$ (9)
Figure 4: Initial state construction, see Appendix A.
@@
Figure 5: Current density DSFs for the honeycomb state presented in Fig. 2 of the main text, with (a) $\mathcal{S}_{j_{x}}$ and (b) $\mathcal{S}_{j_{y}}$. The dotted line in (a) shows the low-$k$ linear dispersion, with gradient equal to the speed of transverse sound.
Appendix B—In order to create panel (c) in Figs. 1 and 2, we take the difference of the two current density DSFs $\mathcal{S}_{j_{x}}$ and $\mathcal{S}_{j_{y}}$. In this Appendix, we wish to highlight the data post-processing required to reduce the noise on this figure, and to find the superfluid, longitudinal crystal, and transverse crystal phonon branches.
In Fig. 5 we show the raw data of the two current density DSFs of Fig. 2(c) on a logarithmic colorscale. Through the numerical noise, the three Goldstone branches are clearly visible. On the transverse branch, we overlay the linear approximation $\omega=c_{t}k_{y}$, showing that the branch visibly exceeds this initially linear trajectory. Other features of interest can be discerned from this figure. The first gapped branch (around $\omega=0.5\omega_{z}$) is brightest in the second BZ, and corresponds to a sub-lattice wavelength shearing of the holes of the honeycomb. A similar feature is shown in (b), where the superfluid phonons of sub-lattice wavelength also couple strongest to this excitation scheme. When plotting Fig. 2 we isolate the peaks of this data, throwing away signals of amplitude $<0.01$ to highlight the relevant branches. We fit these extracted peaks of the three gapless branches with a quartic polynomial for the superfluid and transverse branches, and a tanh curve for the longitudinal branch.
Appendix C—It is pertinent to ask how our results scale in a finite sized system. As the elastic tension for droplets along the edge of the system are only provided by a single neighboring row, can shear waves still propagate in this system? How does the inhomogeneity affect our results? We tackle these questions by considering a system with an additional harmonic trap along one axis, i.e. inserting $V=m\omega_{y}^{2}y^{2}/2$ into our Hamiltonian in Eq. (9), such that our system emulates a segment of a supersolid in a ring trap geometry [89, 90, 91].
[…]
Figure 6 shows our results. The initial state, in (a), is excited via a vertical shift applied to the outermost column of droplets. We find that, as in the infinite supersolid, a shear wave propagates across the system. Unlike the infinite case, there is now a spatially varying velocity, peaking where the density is largest. We define the local velocity between the columns by the time delay at which pairs of columns reach their maximum displacement. This change in velocity is due to the variation in the inter-droplet spacing between columns. Our results verify that shear waves can propagate in an experimentally accessible geometry. In future work, we will investigate this phenomena in two-dimensional box-traps, where the trap itself will play less of a role [92], or in azimuthally symmetric harmonic traps, where the radial shear waves are Tkachenko modes [3, 4, 5, 6].

### v3

@@ Abstract
Dipolar supersolids—quantum states that are simultaneously superfluid and solid—have had their superfluid nature rigorously tested, while its solid nature remains uncharted. Arguably, the defining characteristic of a solid is the existence of elastic shear waves. In this work, we investigate transverse wave packet propagation in dipolar supersolids with triangular and honeycomb structure. Remarkably, the honeycomb supersolid displays anomalous dispersion, supporting waves traveling faster than the transverse speed of sound: a supersonic shear wave. While the distinction between anomalous and normal dispersion appears subtle, it has a profound influence on both the transport and scattering properties. For both supersolid phases, we calculate the shear modulus, a key parameter that quantifies the material’s rigidity. Our findings are pertinent to current experimental efforts scrutinizing the fundamental properties of supersolids.
@@
One other such quantum phase that should support shear waves is the supersolid. A supersolid is a fascinating and counterintuitive state of matter that combines the properties of superfluidity with crystalline order [7, 8, 9, 10, 11, 12]. In this exotic phase, a material exhibits both the rigidity and periodic structure characteristic of an ordered solid, and the inviscid flow typically associated with superfluids. This duality emerges from multiple spontaneously broken symmetries: a global U(1) phase symmetry, and both translation and rotation symmetries [13]. Ultracold atoms have now established themselves as the leading platform to study supersolidity, with varied approaches including cavity settings [14, 15, 16], atoms with spin-orbit coupling [17], and dipolar atoms exhibiting intrinsic long-range interactions [18, 19, 20, 21]. Understanding this phase in the context of ultracold matter may unveil new behavior in solid state materials [22], helium systems [23, 24, 25, 26, 27], and neutron stars [28, 29, 30].
@@
Here, we propose a protocol to excite shear waves in two-dimensional dipolar supersolids, and measure their resulting propagation velocities. We investigate this protocol for triangular and honeycomb lattice structures, which are ground state solutions for a system of ultracold dipolar atoms under planar confinement [41, 42, 43]. Our results reveal distinct dispersive phenomena between the triangular and honeycomb phases, which we probe through the current density dynamic structure factor (DSF) and the Bogoliubov de-Gennes (BdG) method. These investigations show that the honeycomb structure can support a refractive index smaller than unity, thus supporting shear waves traveling faster than the transverse speed of sound. The idea of a supersonic sound mode in classical wave theory can be traced back to Sommerfeld and Brillouin [44, 45, 46, 47], while in the quantum realm, it is analogous to the anomalous phonon dispersion relation observed in superfluid helium-4 [48, 49, 50, 23]. Finally, we find the relationship between the wave packet propagation velocities and the transverse speed of sound, and estimate the shear modulus of dipolar supersolids.
@@
This full system is well described by a beyond mean-field theory, known as the extended Gross-Pitaevskii equation (eGPE) [52, 53, 54, 55]. This governs the three-dimensional macroscopic wavefunction $\Psi(\textbf{r},t)$, and takes into account single-particle dynamics, two-body interactions, and quantum fluctuations. Due to the azimuthal symmetry of the interaction, dipolar supersolids break at most two translational symmetries, giving rise to triangular, stripe, or honeycomb two-dimensional (2D) lattice structures [41, 42, 43]. Thus, following Ref. [42], we are able to reduce the complexity of this model to the quasi-2D variational eGPE
$i\hbar\frac{\partial\psi}{\partial t}=\bigg{[}-\frac{\hbar^{2}\nabla^{2}}{2m}+\Phi_{\sigma}+g_{\text{QF}}|\psi|^{3}\bigg{]}\psi\,,$ (2)
where the dimensionally reduced wave function $\psi(\bm{\rho},t)$ is defined over the infinite plane, $\bm{\rho}=(x,y)$. The 2D interaction and quantum fluctuation terms, $\Phi_{\sigma}$ and $g_{\text{QF}}$, depend on the variational width $\sigma$ in the perpendicular ($z$) direction. We fix the areal number density $n_{\text{2D}}$, and perpendicular harmonic trap with frequency $\omega_{z}=2\pi\times 72.4\mathrm{\,Hz}$. This model and our numerical methods are detailed in the Appendix.
@@
Shear waves in triangular supersolids—In Fig. 1(a), we show the initial perturbation of a single row of the crystal away from its equilibrium position, which adds elastic tension to the lattice. We shift the central row by $\Delta L=0.125L$, where $L$ is the lattice spacing. We have ensured that the observed shear wave propagation does not vary with the magnitude of this shift for smaller perturbations. Evolving the system in real time, we observe a longitudinally propagating wave packet (along $\hat{y}$), leaving behind transverse crystal oscillations (along $\hat{x}$). We then track the deviation of the center of mass (COM) of each row from its equilibrium position, as shown in Fig. 1(b). We track both the peak and edge of this wave packet, defining the latter as a COM deviation larger than 5% of the peak amplitude. The peak velocity increases with the propagation time, starting at $\approx 0.23\,$mm/s and grows to $\approx 0.24\,$mm/s after 750 ms, and the edge velocity always exceeds the peak velocity at $\approx 0.26\,$mm/s, due to the dispersive nature of the wave packet. The existence of shear waves implies a non-zero shear modulus of dipolar supersolids, and proves the solid-like nature of the state, a key result of this work.
@@
Fig. 1(c) shows half of the first Brillouin Zone (BZ), $0<k_{y}<k_{\text{BZ}}=\pi/L$, of the current density DSF, revealing the three gapless (Goldstone) modes inherent to a 2D supersolid [63]. As apparent from the large amplitude positive $\mathcal{S}_{j}$, the response of the system is dominated by $j_{x}$ fluctuations, revealing the transverse branch $\omega_{t}(k_{y})$. Only weak longitudinal crystal and superfluid phonons are induced, as expected from the dislocation in the density at the boundary of the initial perturbation. On the right-hand side of (c), we also highlight the dominant current density fluctuation by taking $\mathcal{F}_{t}[j_{x}(\bm{\rho},t)]=j_{x}(\bm{\rho},\omega)$, and isolating the mode with frequency $\omega_{t}(k_{\text{BZ}})=\omega_{\text{BZ}}$ [64]. We note that $\omega_{\text{BZ}}$ is equal to the oscillation frequency of the central row COM ($y=0$), shown in (b). This mode corresponds to a transverse phonon with wave number $k_{\text{BZ}}$, i.e. a fluctuation where the current alternates between each row.
We independently verify the dispersion of these Goldstone modes by introducing a perturbation about the ground state solution, $\psi(\bm{\rho},t)=\left[\psi_{0}(\bm{\rho})+\delta(\bm{\rho},t)\right]\mathrm{e}^{-\mathrm{i}\mu t/\hbar}$. Here, $\mu$ is the chemical potential and
$\delta(\bm{\rho},t)=\sum_{\nu}\left[u_{\nu}(\bm{\rho})\mathrm{e}^{-\mathrm{i}\omega_{\nu}t}-v_{\nu}^{*}(\bm{\rho})\mathrm{e}^{\mathrm{i}\omega_{\nu}^{*}t}\right],$ (4)
where $u_{\nu}$ and $v_{\nu}$ are the Bogoliubov modes with band index $\nu$ which obey the translational symmetry of the supersolid state. By substituting the perturbed wavefunction into Eq. (2) and linearizing, we obtain the familiar Bogoliubov-de Gennes (BdG) equations [65, 66] (See Appendix C for details). We solve the BdG equations using a Bloch wave expansion of the Bogoliubov modes to obtain the frequencies, $\omega_{\nu}(\bm{k})$, and overlay the three lowest bands for comparison in Fig. 1(c). Note that the near-identical shape of the superfluid and transverse modes is a coincidence of the chosen parameters, and their speeds of sound are distinct in other regions of parameter space [66]. We can extract three characteristic velocities from the BdG results for the transverse branch: 1) the transverse speed of sound $c_{t}=\lim_{k_{y}\rightarrow 0}\omega_{t}/k_{y}$, 2) the phase velocity $v_{p}(k_{y})=\omega_{t}/k_{y}$, which defines the velocity of a wave with given $k_{y}$, and 3) the group velocity, $v_{g}(k_{y})=\partial\omega_{t}/\partial k_{y}$, which defines the velocity of the wave packet envelope and thus the velocity of energy propagation in the medium  [^2] We extract these quantities from the BdG calculations by fitting the transverse branch with a fourth-order polynomial.. For the triangular supersolid, both $v_{g}$ and $v_{p}$ monotonically decrease from $c_{t}$ with increasing wave number, and subsequently the refractive index, $n=c_{t}/v_{p}$, increases from unity with $k_{y}$. This behavior is referred to as normal dispersion and is a hallmark of conventional crystalline solids [68].
@@
We can pinpoint the origin of this unexpected finding from the current density DSF, the BdG dispersion and the corresponding characteristic velocities, as shown in Figs. 2(c)-(d)  [^3] Note that as the perturbation dislocates the high density region, we excite many branches, but still dominantly transverse crystal modes.. The transverse branch displays a subtle upward concavity for intermediate wave vector ($0.2\lesssim k_{y}L/\pi\lesssim 0.4$), supporting supersonic $v_{g}$ and $v_{p}$ ($>c_{t}$), and a sub-unity refractive index. This is known as anomalous dispersion and has significant implications for transverse phonon decay. The superlinear dispersion allows for Beliaev damping of transverse phonons, where one phonon spontaneously decays into two [70, 71]. Though rare, similar phenomena has been observed in other systems such as crystalline materials [72, 73], Wigner and Coulomb crystals [74, 75, 76] and superfluid helium-4 [49], the latter being the subject of thirty years of intense experimental and theoretical debate [23]. We speculate that the physical origin of this anomalous dispersion is related to the recently reported shear instability in the honeycomb supersolid [77], and signals the proximity to a transverse branch instability. We elucidate on this interpretation in the End Matter, however this will be substantiated in a subsequent work. To our knowledge, the honeycomb supersolid constitutes the first quantum solid to display anomalous dispersion, the main result of this Letter.
@@
[…]
Fig. 3 shows the required ingredients in order to estimate the shear modulus from our measured time-dependent velocities $v_{s}$. We generalize our results to a wide-range of 2D densities, whilst maintaining the ground state crystal structure. The shaded region in (a) demarcates the range of velocities that could be observed in an experiment. The lower bound is given by the initial peak velocity (which increases over time), and the upper bound is the edge velocity. Combined with the superfluid fraction calculation of (b), we show that the shear moduli of both states are on the order of attoNewtons per meter (aN/m)  [^5] As the crystal is 2D, the units are in Newtons per length., and are consistent with recent related work [66, 77]. Thus, the small–but non-zero–values of the shear moduli definitively show that the dipolar supersolid phases are solid  [^6] We have verified for both states that these observations are only weakly dependent on the scattering length..
Consider the experimental implementation of this protocol for the triangular supersolid. The edge velocity matches closely with our extracted $c_{t}$ [Fig. 3(a1)], however due to fluctuations such as temperature effects, tracking the edge of the shear wave may be below the fidelity of experiments. Furthermore, in finite size systems the propagation velocity can be greatly reduced at the boundary of the system, as we show in Appendix E. Thus, experimental estimates may drastically underestimate the shear modulus. At much larger densities than shown here, the transverse branch becomes soft, and the shear wave velocity tends to zero. However, this is in the parameter regime where the stripe phase is the ground state solution [42].
@@
Acknowledgements—We acknowledge invaluable conversations with F. Ferlaino, R. Bisset, E. Poli, P. B. Blakie, N. Liebster, M. Olimpo, and the Innsbruck dipolar team. We are particularly grateful to P. B. Blakie for sending us data from Refs. [42] and [66] to benchmark our results. We acknowledge financial support by the ESQ Discovery programme (Erwin Schrödinger Center for Quantum Science & Technology), hosted by the Austrian Academy of Sciences (ÖAW), and the European Research Council through the Advanced Grant DyMETEr (10.3030/101054500). T.B. acknowledges funding from FWF Grant No. I4426.
Appendix C—To validate the dispersion obtained from the current density DSFs, we expand the Bogoliubov modes in Eq. (4) into Bloch waves
$u_{\nu}(\bm{\rho})=u_{\nu,\bm{k}}(\bm{\rho})e^{i\bm{k}\cdot\bm{\rho}},\quad v_{\nu}(\bm{\rho})=v_{\nu,\bm{k}}(\bm{\rho})e^{i\bm{k}\cdot\bm{\rho}}.$ (12)
By substituting the perturbation into the quasi-2D eGPE and linearizing, we obtain the BdG matrix
$\hbar\omega_{\nu}(\bm{k})\begin{bmatrix}u_{\nu,\bm{k}}\\ v_{\nu,\bm{k}}\end{bmatrix}=\begin{bmatrix}\mathcal{L}+\mathcal{M}&-\mathcal{M}\\ \mathcal{M}&-\left(\mathcal{L}+\mathcal{M}\right)\end{bmatrix}\begin{bmatrix}u_{\nu,\bm{k}}\\ v_{\nu,\bm{k}}\end{bmatrix},$ (13)
where the operators are
$\mathcal{L}$ $=-\frac{\hbar^{2}\nabla^{2}}{2m}-\mu+\Phi_{\sigma}+g_{\text{QF}}|\psi|^{3},$ (14)
$\mathcal{M}f$ $=\psi_{0}\mathcal{F}^{-1}\left[\tilde{U}(\bm{k})\mathcal{F}\left[f\psi_{0}^{*}\right]\right]+\frac{3}{2}g_{\mathrm{QF}}|\psi_{0}|^{3}f.$ (15)
We fit the transverse branch obtained from this procedure with a quartic polynomial in order to extract the three characteristic velocities shown in panel (d) in Figs. 1 and 2. In Fig. 6 we show the transverse branches and the quartic fit for the triangular and honeycomb supersolids at $n_{\text{2D}}=2000\,\mu$m-2 and $n_{\text{2D}}=4300\,\mu$m-2 respectively. Similar to Fig. 5, we overlay the linear approximation $\omega=c_{t}k_{y}$, showing that the transverse branch of the honeycomb supersolid exceeds this linear trajectory in Fig. 6(b).
Appendix D—Here we present a simple physical argument for the origin of the anomalous behavior, motivated by the recent work on the shear instability in the honeycomb supersolid of Ref. [77]. In general, the dispersion relation for a gapless branch must have the form $\omega(k)=Ak+Bk^{2}+\mathcal{O}(k^{3})$ for momentum $k$. A phonon instability occurs when the linear part becomes negative, crossing $A=0$. However, in a crystalline system at the Brillouin zone (BZ) edge $\text{d}\omega(k)/\text{d}k=0$. Therefore, near to the shear wave (transverse phonon) instability the band goes as $k^{2}$ for low momenta and flattens at the BZ edge. These criteria set “boundary conditions” for the band and lead to anomalous behavior for intermediate $k$. We note that this implies an enhancement of the anomalous behavior close to the shear wave instability, however there the honeycomb state is not the ground state, and it may not be possible to dynamically probe this in an experiment without entering the stripe regime. Our results show that this behavior persists in an experimentally relevant parameter regime where the honeycomb state is indeed the ground state.
Appendix E—It is pertinent to ask how our results scale in a finite sized system. As the elastic tension for droplets along the edge of the system are only provided by a single neighboring row, can shear waves still propagate in this system? How does the inhomogeneity affect our results? We tackle these questions by considering a system with an additional harmonic trap along one axis, i.e. inserting $V=m\omega_{y}^{2}y^{2}/2$ into our Hamiltonian in Eq. (2), such that our system emulates a segment of a supersolid in a ring trap geometry [99, 100, 101].
@@
Fig. 7 shows our results. The initial state, in (a), is excited via a vertical shift applied to the outermost column of droplets. We find that, as in the infinite supersolid, a shear wave propagates across the system. Unlike the infinite case, there is now a spatially varying velocity, peaking where the density is largest. We define the local velocity between the columns by the time delay at which pairs of columns reach their maximum displacement. This change in velocity is due to the variation in the inter-droplet spacing between columns. Our results verify that shear waves can propagate in an experimentally accessible geometry. In future work, we will investigate this phenomena in two-dimensional box-traps, where the trap itself will play less of a role [102], or in azimuthally symmetric harmonic traps, where the radial shear waves are Tkachenko modes [3, 4, 5, 6].